# Improving Medium Range Severe Weather Prediction through Transformer Post-processing of AI Weather Forecasts

### v2

@@ Abstract
Improving the skill of medium-range (1-8 day) severe weather prediction is crucial for mitigating societal impacts. This study introduces a novel approach leveraging decoder-only transformer networks to post-process AI-based weather forecasts, specifically from the Pangu-Weather model, for improved severe weather guidance. Unlike traditional post-processing methods that use a dense neural network to predict the probability of severe weather using discrete forecast samples, our method treats forecast lead times as sequential “tokens”, enabling the transformer to learn complex temporal relationships within the evolving atmospheric state. We compare this approach against post-processing of the Global Forecast System (GFS) using both a traditional dense neural network and our transformer, as well as configurations that exclude convective parameters to fairly evaluate the impact of using the Pangu-Weather AI model. Results demonstrate that the transformer-based post-processing significantly enhances forecast skill compared to dense neural networks. Furthermore, AI-driven forecasts, particularly Pangu-Weather initialized from high resolution analysis, exhibit superior performance to GFS in the medium-range, even without explicit convective parameters. Our approach offers improved accuracy, and reliability, which also provides interpretability through feature attribution analysis, advancing medium-range severe weather prediction capabilities.
@@ 1 Introduction
Traditionally, severe weather forecasting has relied heavily on Numerical Weather Prediction (NWP) models, which provide both environmental and storm-scale information that can be used to infer the likelihood of convective hazards in the future. This includes the use of convection-allowing models (CAMs) at lead-times of 1–2 days (e.g., the High-Resolution Rapid Refresh model (Dowell et al., 2022)) and coarser NWP ensembles at lead times of up to and beyond a week (e.g., the Global Ensemble Forecasting System). Since these NWP systems do not explicitly represent convective hazards, machine learning (ML) can be used to post-process NWP output, using predicted environmental and storm-scale fields to generate hazard probabilities (McGovern et al., 2023). ML models such as random forests (Hill et al., 2020, 2023; Loken et al., 2020, 2022), dense neural networks (Sobash et al., 2020; Sobash and Ahijevych, 2024, 2025), and convolutional neural networks (Gagne et al., 2019; Sha et al., 2024) have all been used to generate skillful guidance for convective hazards over CONUS using operational NWP model output at lead-times from hours to days.
@@
This study extends this line of research by introducing and evaluating two novel approaches to severe weather prediction: 1) the use of AI NWP emulators to provide forecasts of environmental conditions, and 2) a post-processing system to generate convective hazard probabilities based on the use of transformer networks, rather than random forests or dense neural networks as in prior work. AI NWP emulators, like Pangu-Weather (Bi et al., 2023), offer a significant advantage in terms of computational speed and have demonstrated comparable or even superior performance to operational global NWP forecasts (Rasp et al., 2023). A recent study by Hua et al. (2025) shows that Pangu-Weather is more accurate than GFS in forecasting wind shear and storm-relative helicity at the grid point and hour closest to the tornado report at a one-day lead time. We hypothesize that gains in large-scale forecast skill will translate into more skillful convective hazard forecasts, especially in the medium-range (i.e., lead-times of 3–8 days) where the skill differences between traditional and data-driven NWP are most significant.
@@
To generate hazard probabilities, we post-process the AI NWP forecast output using decoder-only transformers. To our knowledge, this is the first study to use transformers to make severe weather predictions. Taking inspiration from large language modeling, we treat each forecast lead time as a "token" within a sequence. Rather than predicting the next token in the sequence, our model predicts the probability of severe weather for each day. The transformer architecture allows the model to learn complex temporal relationships and dependencies within the sequence of input forecasts, effectively capturing the evolution of atmospheric conditions leading to severe weather.
@@
In this work, we document a set of experiments to test the impact of the use of AI NWP emulators and transformers for convective hazard prediction by comparing convective hazard forecast skill to existing baselines during 2024, an active severe weather season across CONUS. Hazard forecasts are generated using Pangu-Weather initialized with several different initial conditions and are compared to forecasts using the Global Forecast System (GFS). In addition, hazard forecasts generated with a transfomer-based post-processing system are compared to using dense neural networks. We also explore combining the data-driven AI and traditional NWP forecasts to better capture forecast uncertainty in an ensemble approach. Together, these experiments provide insight into the ability of AI NWP emulators to improve forecasts of small-scale hazards and extend practical predictability of convective hazards into the medium-range.
@@ 2.1 Preprocessing Global NWP and AI Forecast Dataset
The Pangu-Weather AI NWP model emulator (Bi et al., 2023) leverages a 3D Earth-specific transformer with hierarchical temporal aggregation for fast and accurate global weather forecasting. Trained on 1979–2017 ERA5 (Hersbach et al., 2020) reanalysis data with a resolution of 0.25 by 0.25-degree, four separate networks predict mean sea level pressure, surface temperature, wind, and 5 pressure-level variables at lead times of 1, 3, 6, and 24 hours. The 5 pressure-level variables are geopotential height, temperature, specific humidity, and U and V components of wind at 13 pressure levels. The set of variables used as input to the model is identical to the set of variables it predicts as output, enabling iterative prediction. Pangu-Weather employs a hierarchical temporal aggregation approach, which is a greedy algorithm designed to forecast in the fewest possible steps. For example, to generate a 54-hour lead time forecast, Pangu-Weather iteratively applies the 24-hour model twice, followed by the 6-hour model. We generate and save 6-hourly output for all Pangu-Weather forecasts.
We use Pangu-Weather to generate daily 00 UTC-initialized forecasts from 1 January 2018 – 1 August 2024 with three different initial conditions (ICs): the ERA5 reanalysis data, the ECMWF High-resolution (HRES) analysis (used to initialize the ECMWF Integrated Forecasting System), and GFS analysis. The HRES analysis has a nominal grid point spacing of 9 km and was bilinearly interpolated to match the grid point spacing of ERA5 reanalysis using xESMF (Zhuang, 2019). It is important to note that Pangu-Weather was trained with ERA5 and not been fine-tuned on either HRES or GFS analyses. Using HRES and GFS for ICs allows us to assess its potential performance using operational datasets, since ERA5 reanalysis data is not available in real-time.
@@ 5 Labeling: Binary label for each grid point and lead time range:
The gridded forecast output was upscaled onto an 80-km grid covering all of CONUS, similar to the pre-processing applied in (Sobash et al., 2020). Here, we use inverse distance weighting (IDW) to estimate values at 80-km grid points. The IDW weights are determined by a Gaussian kernel with a 25-km standard deviation, applied within a 120-km search radius. The model variables and dataset preprocessing details are provided in Table 1 and Table 2, respectively. Instead of using all available forecast variables, a subset of key variables known to indicate favorable synoptic conditions for severe weather is selected as input. For each 80-km grid point across the CONUS, and for daily initialization times at 00 UTC, forecast data is aggregated for eight lead time ranges (days 1–8, 12 UTC – 12 UTC). For every 6-hourly time step in each 24-hour lead time range, the grid point’s variables’ value are extracted and also the variables’ mean value are calculated from a 5x5 grid neighborhood centered around the nearest grid point.
The distances to severe weather events from the NOAA Storm Prediction Center (SPC) archive were recorded (but not used as a predictive feature), and a binary label is assigned to each grid point indicating whether a severe report occurred within a 40-km radius during the corresponding 24-hour period. The normalization constants, mean and standard deviation, were calculated using all forecasts from 1 January 2018–31 December 2023. The construction of the preprocessed normalized dataset varies depending on the post-processing framework used, as described in the following Section 2.2.
@@ 2.2 Post-processing Forecast with Decoder-only Transformer
Previous work by Sobash et al. (2020) used a dense neural network (DNN) for post-processing, as shown in Fig. 1a. The DNN takes a vector of discrete forecast samples and predicts the probability of severe weather events for a given location and lead-time. For each forecast day and grid point, the features include those listed in Table 1 at each 6-hourly output time within the 24-hour forecast period. Instead of using features valid within a specific 24-hour period as input to a DNN, the decoder-only transformer network (DOT; Vaswani 2017, Radford et al. 2019) uses forecast information across all forecast days, as shown in Fig. 1b. The input is a matrix instead of a vector, with rows representing lead time and columns representing features, and the output is a sequence of severe weather probabilities for each leadtime. This mimics how tokens in large language modeling are structured, where a sentence is broken down into a series of tokens (i.e., words) stacked over the row dimension, and each token has a certain embedding represented in the column dimension.
@@
Compared to the DNN, using the DOT in this way allows the model to use forecast information at multiple days for a given forecast point, and the attention mechanism learns how to relate variables to one another. The specific architecture uses a ‘TransformerDecoder‘ with several key hyperparameters detailed in Fig. 1b. The ‘d_model’ parameter defines the dimensionality of the embeddings and internal layers, essentially setting the capacity for representing features at each time step. The model consists of ‘num_layers’ identical decoder layers stacked sequentially, allowing for hierarchical processing of the input sequence. Within each layer, the multi-head attention mechanism uses ‘nhead’ parallel attention heads, enabling the model to simultaneously focus on different types of relationships between forecast variables across different lead times when predicting the severe weather probability. Additionally, the attention mechanism incorporates a crucial look-ahead mask, enabled by ‘tgt_mask=True’. This mask enforces the auto-regressive property, ensuring that when predicting the severe weather probability for a specific lead time (e.g., Day 4), the model can only attend to the forecast information from the current and preceding lead times (Days 1-4) and is prevented from accessing information from future lead times (Days 5-8). This is analogous to sentence generation where the model cannot look ahead at future tokens when predicting the current token. During inference, we maintain ‘tgt_mask=True’; the model receives the consecutive forecast sample from Day 1 to Day 8 as input and simultaneously outputs the severe weather probability for each day in the sequence (Day 1 to Day 8), from which we extract the desired probability for each day.
@@
Figure 1: Comparison of post-processing frameworks: (a) The Dense Neural Network (DNN) framework takes a discrete sample with an input shape of (Batch, Feature) and outputs a severe weather probability for a single day, with an output shape of (Batch, 1). Shapes above arrows indicate the output shape of each layer. (b) The Decoder-only Transformer (DOT) framework utilizes a consecutive sample of weather forecasts from Day 1 to Day 8, with an input shape of (Batch, Sequence, Feature), and outputs severe weather probabilities for each day from Day 1 to Day 8, resulting in an output shape of (Batch, Sequence, 1). Description of hyperparameters for the PyTorch (Paszke, 2019) TransformerDecoder (‘d_model‘: embedding/layer dimension, ‘num_layers‘: number of stacked layers, ‘nhead‘: number of attention heads, ‘tgt_mask=True‘: enables causal masking) are further elaborate in Section 2.2.
@@ 2.3.1 Model and Training Setup
Each model was trained on forecast data from January 1, 2018, to December 31, 2023. The validation set excluded the period between May 15, 2023, and August 15, 2023. The test period forecasts were initialized using data from February 1, 2024, to July 22, 2024, and with forecast valid times extending to July 31, 2024. This short period exclusion in the validation set was a minor adjustment to maximize the training dataset, as validation was primarily used for early stopping. The training objective is identical to that of a binary classification task, which uses binary cross-entropy loss. A batch size of 512 was consistently used for both the transformer and DNN frameworks. Both frameworks utilized a Cosine Annealing Warm Restarts scheduler. This scheduler cyclically adjusts the learning rate, starting with an initial period of 10 epochs, and subsequently maintaining the period between restarts after each cycle, with a minimum learning rate of 1e-7. The optimizer for both was Stochastic Gradient Descent with a momentum of 0.9 and a weight decay of 1e-4. The initial learning rate for the transformer was set to 6e-4, while for the dense neural network it was 8e-4. Training was halted if no improvement in the validation Brier Skill Score was observed for 20 consecutive epochs. Validation BSS typically plateaued after approximately 150 epochs for the transformer model and around 50 epochs for the dense neural network model. Finally, the training was performed on a single Nvidia L40 GPU.

### v3

@@ Abstract
Improving the skill of medium-range (3-8 day) severe weather prediction is crucial for mitigating societal impacts. This study introduces a novel approach leveraging decoder-only transformer networks to post-process AI-based weather forecasts, specifically from the Pangu-Weather model, for improved severe weather guidance. Unlike traditional post-processing methods that use a dense neural network to predict the probability of severe weather using discrete forecast samples, our method treats forecast lead times as sequential “tokens”, enabling the transformer to learn complex temporal relationships within the evolving atmospheric state. We compare this approach against post-processing of the Global Forecast System (GFS) using both a traditional dense neural network and our transformer, as well as configurations that exclude convective parameters to fairly evaluate the impact of using the Pangu-Weather AI model. Results demonstrate that the transformer-based post-processing significantly enhances forecast skill compared to dense neural networks. Furthermore, AI-driven forecasts, particularly Pangu-Weather initialized from high resolution analysis, exhibit superior performance to GFS in the medium-range, even without explicit convective parameters. Our approach offers improved accuracy, and reliability, which also provides interpretability through feature attribution analysis, advancing medium-range severe weather prediction capabilities.
@@ 1 Introduction
Traditionally, severe weather forecasting has relied heavily on Numerical Weather Prediction (NWP) models, which provide both environmental and storm-scale information that can be used to infer the likelihood of convective hazards in the future. This includes the use of convection-allowing models (CAMs) at lead-times of 1–2 days (e.g., the High-Resolution Rapid Refresh model (Dowell et al., 2022)) and coarser NWP ensembles at lead times of up to and beyond a week (e.g., the Global Ensemble Forecasting System). Since these NWP systems do not explicitly represent convective hazards, machine learning (ML) can be used to post-process NWP output (Vannitsem et al., 2021), using predicted environmental and storm-scale fields to generate hazard probabilities (McGovern et al., 2023). ML models such as random forests (Hill et al., 2020, 2023; Loken et al., 2020, 2022), dense neural networks (Sobash et al., 2020; Sobash and Ahijevych, 2024, 2025), and convolutional neural networks (Gagne et al., 2019; Sha et al., 2024) have all been used to generate skillful guidance for convective hazards over CONUS using operational NWP model output at lead-times from hours to days. Beyond convective hazards, several works have focused on postprocessing forecasts for extreme events such as high-wind scenarios (Veldkamp et al., 2021; Schulz and Lerch, 2022).
@@
This study extends this line of research by introducing and evaluating two novel approaches to severe weather prediction: 1) a post-processing system to generate convective hazard probabilities based on the use of transformer networks, rather than random forests or dense neural networks as in prior work, and 2) the use of AI NWP emulators to provide forecasts of environmental conditions. AI NWP emulators, like Pangu-Weather (Bi et al., 2023), offer a significant advantage in terms of computational speed and have demonstrated comparable or even superior performance to operational global NWP forecasts (Rasp et al., 2023). Recent work has also begun to explore the post-processing of AI-based weather model output for various applications, such as ensemble postprocessing for uncertainty quantification (Bülte et al., 2025) and forecast refinement (Bremnes et al., 2024). These studies also test the ability of such models to reliably simulate other hazards like tropical cyclones (DeMaria et al., 2024), heatwaves (Meng et al., 2025), and to reconstruct derecho climatology (Li et al., 2024) from postprocessed reanalysis and observations. A recent study by Hua et al. (2025) shows that Pangu-Weather is more accurate than GFS in forecasting wind shear and storm-relative helicity at the grid point and hour closest to the tornado report at a one-day lead time. We hypothesize that gains in large-scale forecast skill will translate into more skillful convective hazard forecasts, especially in the medium-range (i.e., lead-times of 3–8 days) where the skill differences between traditional and data-driven NWP are most significant.
@@
To generate hazard probabilities, we post-process the AI NWP forecast output using decoder-only transformers. To our knowledge, this is the first study to use transformers for severe weather prediction, although other recent papers have proposed transformer-based post-processing models (Bouallègue et al., 2024; Van Poecke et al., 2025) applied to forecast fields. Taking inspiration from large language modeling, we treat each forecast lead time as a "token" within a sequence. Rather than predicting the next token in the sequence, our model predicts the probability of at least one severe weather report (any hazard) within a 40 km radius for 24-hour periods from day 1 through day 8. The transformer architecture allows the model to learn complex temporal relationships and dependencies within the sequence of input forecasts, effectively capturing the evolution of atmospheric conditions leading to severe weather.
@@
In this work, we document a set of experiments to test the impact of the use of AI NWP emulators and transformers for convective hazard prediction by comparing convective hazard forecast skill to existing baselines during 2024, an active severe weather season across CONUS. Weather forecasts are generated using Pangu-Weather initialized with several different initial conditions and are compared to forecasts from the Global Forecast System (GFS). In addition, the forecasts are post-processed to generate hazard forecasts, and the transformer-based post-processing system is compared against dense neural networks. We also explore combining the data-driven AI and traditional NWP forecasts to better capture forecast uncertainty in an ensemble approach. Together, these experiments provide insight into the ability of AI NWP emulators to improve forecasts of small-scale hazards and extend practical predictability of convective hazards into the medium-range.
@@ 2.1 Preprocessing Global NWP and AI Forecast Dataset
The Pangu-Weather AI NWP model emulator (Bi et al., 2023) leverages a 3D Earth-specific transformer with hierarchical temporal aggregation for fast and accurate global weather forecasting. Trained on 1979–2017 ERA5 (Hersbach et al., 2020) reanalysis data with a resolution of 0.25 by 0.25-degree, four separate networks are used to predict at four different lead times (1, 3, 6, and 24 hours) for mean sea level pressure, surface temperature, wind, and 5 pressure-level variables. The 5 pressure-level variables are geopotential height, temperature, specific humidity, and U and V components of wind at 13 pressure levels. The set of variables used as input to the model is identical to the set of variables it predicts as output, enabling iterative prediction. Pangu-Weather employs a hierarchical temporal aggregation approach, which is a greedy algorithm designed to forecast in the fewest possible steps. For example, to generate a 54-hour lead time forecast, Pangu-Weather iteratively applies the 24-hour model twice, followed by the 6-hour model. We generate and save 6-hourly output for all Pangu-Weather forecasts.
We use Pangu-Weather to generate daily 00 UTC-initialized forecasts from 1 January 2018 – 1 August 2024 with three different initial conditions (ICs): the ERA5 reanalysis data, the ECMWF High-resolution (HRES) analysis (used to initialize the ECMWF Integrated Forecasting System), and GFS analysis. The HRES analysis has a nominal grid point spacing of 9 km and was bilinearly interpolated to match the grid point spacing of ERA5 reanalysis using xESMF (Zhuang, 2019). It is important to note that Pangu-Weather was trained with ERA5 and has not been fine-tuned on either HRES or GFS analyses. Using HRES and GFS for ICs allows us to assess its potential performance using operational datasets, since ERA5 reanalysis data is not available in real-time.
@@ 5 Labeling: Binary label for each grid point and lead time range:
The model variables and dataset preprocessing details are provided in Table 1. Instead of using all available forecast variables, a subset of key variables known to indicate favorable synoptic conditions for severe weather is selected as input. Table 1 presents the forecast variables used in the experiment. Specifically, “GFS” contains the same variables as the “Pangu” variants, with the different Pangu variants representing different initial conditions. “GFS+Convect” includes several convective parameters and surface variables available from the GFS model output which are not available in the Pangu-Weather model output. This allows for a comparison of the importance of convective parameters in severe weather forecasting, especially since most AI weather models lack these parameters.
The preprocessing details are provided in Table 2. The gridded forecast output was upscaled onto an 80-km grid covering all of CONUS, similar to the pre-processing applied in (Sobash et al., 2020). Here, we use inverse distance weighting (IDW) to estimate values at 80-km grid points. The IDW weights are determined by a Gaussian kernel with a 25-km standard deviation, applied within a 120-km search radius. For each 80-km grid point across the CONUS, and for daily initialization times at 00 UTC, forecast data is aggregated for eight lead time ranges (days 1–8, 12 UTC – 12 UTC). For every 6-hourly time step in each 24-hour lead time range, the output variables at each grid point are extracted. Additionally, the mean of these variables over a 5x5 neighborhood centered on the grid point is calculated. Similar to findings in Hill et al. (2023), adding additional statistics such as “min” and “max” or increasing the neighborhood size showed no significant improvement in model performance.
The distances to severe weather events from the NOAA Storm Prediction Center (SPC) archive were recorded, and a binary label is assigned to each grid point indicating whether a severe report occurred within a 40-km radius during the corresponding 24-hour period. While this distance information was not used as a predictive feature, it was recorded to allow for flexibility in adjusting the proximity threshold for other potential applications. The normalization constants, mean and standard deviation, were calculated using all forecasts from 1 January 2018–31 December 2023. The construction of the preprocessed normalized dataset varies depending on the post-processing framework used, as described in the following Section 2.2.
@@ 2.2 Post-processing Forecast with Decoder-only Transformer
Previous work by Sobash et al. (2020) used a dense neural network (DNN) for post-processing, as shown in Fig. 1a. The DNN takes a vector of discrete forecast samples and predicts the probability of severe weather events for a given location and lead-time. For each forecast day and grid point, the features include those listed in Table 1 at each 6-hourly output time within the 24-hour forecast period. Instead of using features valid within a specific 24-hour period as input to a DNN, the decoder-only transformer network (DOT; Vaswani 2017, Radford et al. 2019) uses forecast information across all forecast days, as shown in Fig. 1b. The input is a matrix instead of a vector, with rows representing lead time and columns representing features, and the output is a sequence of severe weather probabilities for each leadtime at a single grid point. This mimics how tokens in large language modeling are structured, where a sentence is broken down into a series of tokens (i.e., words) stacked over the row dimension, and each token has a certain embedding represented in the column dimension.
@@
Compared to the DNN, using the DOT in this way allows the model to use forecast information at multiple days for a given forecast point, and the attention mechanism learns how to relate variables to one another. The specific architecture uses a ‘TransformerDecoder‘ with several key hyperparameters detailed in Fig. 1b. The ‘d_model’ parameter defines the dimensionality of the embeddings and internal layers, essentially setting the capacity for representing features at each time step. The model consists of ‘num_layers’ identical decoder layers stacked sequentially, allowing for hierarchical processing of the input sequence. Within each layer, the multi-head attention mechanism uses ‘nhead’ parallel attention heads, enabling the model to simultaneously focus on different types of relationships between forecast variables across different lead times when predicting the severe weather probability. Additionally, the attention mechanism incorporates a crucial look-ahead mask, enabled by ‘tgt_mask=True’. This mask enforces the auto-regressive property and is used both in training and inference, ensuring that when predicting the severe weather probability for a specific lead time (e.g., Day 4), the model can only attend to the forecast information from the current and preceding lead times (Days 1-4) and is prevented from accessing information from future lead times (Days 5-8). This auto-regressive property is enforced to ensure the model learns the temporal evolution of atmospheric features leading up to an event, preventing reliance on forecast data from future lead times which could obscure causal relationships. This is analogous to sentence generation where the model cannot look ahead at future tokens when predicting the current token. During inference, we maintain ‘tgt_mask=True’; the model receives the consecutive forecast sample from Day 1 to Day 8 as input and simultaneously outputs the severe weather probability for each day in the sequence (Day 1 to Day 8), from which we extract the desired probability for each day.
@@
Figure 1: Comparison of post-processing frameworks: (a) The Dense Neural Network (DNN) framework takes a discrete sample with an input shape of (Batch, Feature) and outputs a severe weather probability for a single day, with an output shape of (Batch, 1). Shapes above arrows indicate the output shape of each layer. (b) The Decoder-only Transformer (DOT) framework utilizes a consecutive sample of weather forecasts from Day 1 to Day 8, with an input shape of (Batch, Sequence, Feature), and outputs severe weather probabilities for each day from Day 1 to Day 8, resulting in an output shape of (Batch, Sequence, 1). Description of hyperparameters for the PyTorch (Paszke, 2019) TransformerDecoder (“d_model”: embedding/layer dimension, “num_layers”: number of stacked layers, “nhead”: number of attention heads, “tgt_mask=True”: enables causal masking) are further elaborate in Section 2.2.
@@ 2.3.1 Model and Training Setup
Each model was trained on forecast data from January 1, 2018, to December 31, 2023. The validation set excluded the period between May 15, 2023, and August 15, 2023. The test period forecasts were initialized using data from February 1, 2024, to July 22, 2024, and with forecast valid times extending to July 31, 2024. This short period exclusion in the validation set was a minor adjustment to maximize the training dataset, as validation was primarily used for early stopping. The training objective is identical to that of a binary classification task, which uses binary cross-entropy loss. A batch size of 512 was consistently used for both the transformer and DNN frameworks. Both frameworks utilized a Cosine Annealing Warm Restarts scheduler. This scheduler cyclically adjusts the learning rate, starting with an initial period of 10 epochs, and subsequently maintaining the period between restarts after each cycle, with a minimum learning rate of 1e-7. The optimizer for both was Stochastic Gradient Descent with a momentum of 0.9 and a weight decay of 1e-4. The initial learning rate for the transformer was set to 6e-4, while for the dense neural network it was 8e-4. Training was halted if no improvement in the validation Brier Skill Score was observed for 20 consecutive epochs. Validation BSS typically plateaued after approximately 150 epochs for the transformer model and around 50 epochs for the dense neural network model. Finally, the training was performed on a single Nvidia L40 GPU. Additional ablation tests regarding “d_model” and “num_layers” are provided in Fig. S1 and show that there was no significant impact on performance when the other settings remained the same.